# Non-Gravitational Acceleration in 3I/ATLAS: Constraints on Exotic Volatile Outgassing in Interstellar Comets

### v2

@@ Abstract
comets: general – comets: individual (3I/ATLAS) – interstellar objects – methods: numerical – celestial mechanics
pubyear: 2025pagerange: Non-Gravitational Acceleration in 3I/ATLAS: Constraints on Exotic Volatile Outgassing in Interstellar CometsC
@@ 2 Observational Constraints
A non-gravitational solution was retrieved from HORIZONS on 15 January 2026. This solution reports statistically significant radial and transverse components, expressed in units of AU day-2 and converted to SI units using $1~\mathrm{AU~day^{-2}}=2.004\times 10^{-4}~\mathrm{m~s^{-2}}$. The resulting values are
@@
These correspond to a net non-gravitational acceleration of approximately
$a_{\mathrm{ng}}=(5\pm 2)\times 10^{-7}~\mathrm{m~s^{-2}}$
@@
To verify that the inclusion of non-gravitational terms improves the orbital fit, we compared the residuals of the gravitational-only and full HORIZONS solutions. The gravitational-only solution exhibits systematic residual trends aligned primarily with the heliocentric radial direction. In contrast, the solution including non-gravitational parameters removes these trends and reduces the residuals to the level of the astrometric uncertainties. This behavior is characteristic of recoil forces arising from anisotropic mass loss and is consistent with expectations for an active cometary body.
The magnitude of the inferred acceleration lies well within the range observed for modestly active comets in the inner Solar System. Typical non-gravitational accelerations for comets near 1 AU span $10^{-9}$ to $10^{-6}~\mathrm{m~s^{-2}}$, depending on nucleus size and volatile composition (Yeomans2004; Combi2011). For comparison, the interstellar object 1I/‘Oumuamua exhibited an acceleration of order $3\times 10^{-6}~\mathrm{m~s^{-2}}$ (Micheli2018). The value inferred for 3I/ATLAS is smaller by roughly an order of magnitude, consistent with a weaker level of activity and compatible with standard volatile-driven outgassing at heliocentric distances of order 1 to 2 AU.
@@ 3.3 Material Parameters and Model Inputs
The physical parameters adopted in the model are summarized in Table 1. Molecular masses and latent heats are taken from laboratory measurements, while albedos and emissivities are representative of low-albedo cometary nuclei. The density is fixed at $500\,\mathrm{kg\,m^{-3}}$, consistent with in situ measurements of cometary material (Patzold2016). The diurnal illumination factor $\mu_{0}$ depends on both rotation period and obliquity; because these are not constrained for 3I/ATLAS, $\mu_{0}$ is computed using a diurnal average for a range of assumed spin-axis orientations.
These inputs determine the equilibrium surface temperature, mass flux, and resulting acceleration for each volatile species at different heliocentric distances.
Table 1: Thermophysical and compositional parameters adopted for modeling the sublimation of CO, CO2, NH3, and CH4.
Parameter CO CO2 NH3 CH4
Latent heat $L_{i}$ (J kg-1) $2.7{\times}10^{5}$ $5.7{\times}10^{5}$ $1.4{\times}10^{6}$ $5.1{\times}10^{5}$
Molecular mass $\mu_{i}$ (kg mol-1) 0.028 0.044 0.017 0.016
Bond albedo AA 0.04 0.04 0.04 0.04
Emissivity $\varepsilon$ 0.9 0.9 0.9 0.9
Figure 1 illustrates the modelled non-gravitational acceleration as a function of heliocentric distance for each volatile. The calculations use diurnal-averaged illumination and a representative collimation factor $\eta=2.5$. Under these assumptions, CO provides by far the strongest recoil acceleration over the range $0.4\lesssim r_{H}\lesssim 3$ AU, NH3 and CH4 yield weaker but non-negligible accelerations closer to the Sun, and CO2 produces effectively no recoil because its vapor pressure remains extremely low at the corresponding temperatures.
Figure 1: Modeled non-gravitational acceleration as a function of heliocentric distance for CO, CO2, NH3, and CH4. Accelerations are computed from the thermophysical model using diurnal-averaged insolation, empirical vapor-pressure relations, and a representative jet-collimation factor $\eta=2.5$. The example calculation shown adopts $R_{\mathrm{nuc}}=5.6$ km, bulk density $\rho=500$ kg m-3, and total active fraction $f_{\mathrm{act}}=0.02$. The curves illustrate the relative efficiency of each volatile species at heliocentric distances relevant to the 3I/ATLAS observations.
@@ 4.3 Comparison with Observations
Because the direction of the non-gravitational acceleration is poorly constrained for 3I/ATLAS, we match only the magnitude of the modeled acceleration to the scalar value $|\vec{a}_{\mathrm{obs}}|$ inferred from the Marsden $A_{1}$ and $A_{2}$ terms. For each nucleus radius and volatile configuration, we determine the minimum total active fraction $f_{\mathrm{act,tot}}$ required for
@@
This scalar-matching procedure produces a well-defined lower bound on the required active area, independent of directional degeneracies that cannot be resolved from existing data.
Figure 2 shows the resulting required active fractions for nucleus radii between 0.5 and 3 km. The single-species curves (CO, CO2, NH3) rise from $f_{\mathrm{act}}\sim 0.1$ to order unity as radius increases, reflecting the fact that sublimation-driven recoil scales as active area divided by nucleus mass. At the adopted heliocentric distance of $r_{H}=1.36$ AU, pure CO and CO2 require large active areas to match the observed acceleration, consistent with their moderate vapor pressures at these temperatures.
The CO–CO2 mixture curve yields extremely small scalar active fractions. This efficiency does not correspond to a physically unique solution; instead, it reflects an idealized mathematical lower bound obtained by combining species contributions without enforcing directional consistency. Any physically realizable configuration that also matches the acceleration’s vector direction would require equal or larger total active area. The mixture curve should therefore be interpreted only as a formal lower limit rather than as a realistic description of the surface activity of 3I/ATLAS.
Figure 2: Required total active fraction as a function of nucleus radius for single-volatile and mixed-volatile models. For each radius, the scalar acceleration $|\vec{a}_{\mathrm{ng}}|$ produced by the modeled outgassing is matched to the observed magnitude at $r_{H}=1.36$ AU. Single-species curves (CO, CO2, NH3) show the expected increase in required active area with nucleus size, reflecting the mass-to-area scaling of sublimation-driven recoil. The CO–CO2 mixture curve represents an idealized upper limit obtained by scalar combination of species contributions and should not be interpreted as a unique physical solution; directional fits would yield larger required active fractions.
@@ 5.4 Acceleration and Thermophysical Profiles
Figure 1 shows the modeled non-gravitational acceleration as a function of $r_{H}$. CO produces the dominant recoil. NH3 and CH4 produce weaker, but still non-negligible, accelerations closer to the Sun. CO2 remains ineffective over this entire range.
Figure 2 shows that smaller nuclei require smaller active fractions due to the mass-to-area scaling. Mixed CO–CO2 results represent mathematical minima and not physically unique or directional solutions.
Figure 3 shows the latent-to-radiative power ratio $\Lambda(r_{H})$:
  • $\Lambda\gg 1$ (CO, NH3, CH4): sublimation-dominated regime.
  • $\Lambda\ll 1$ (CO2): radiative-dominated regime; negligible recoil.
This hierarchy reinforces that CO must dominate any recoil-producing activity.
Figure 3: Latent-to-radiative power ratio $\Lambda=(ZL)/(\varepsilon\sigma T^{4})$ as a function of heliocentric distance for CO, CO2, NH3, and CH4, using diurnal-averaged illumination. $\Lambda\gg 1$ corresponds to sublimation-dominated surface energy balance, while $\Lambda\ll 1$ indicates radiative dominance. CO, NH3, and CH4 remain sublimation-dominated across 0.4–3 AU, whereas CO2 remains radiatively dominated and therefore cannot generate significant recoil.
@@ 7 Conclusions
Acknowledgements
The authors thank the Leiden Institute of Advanced Computer Science at Leiden University for institutional support and access to computational resources used in this work. The authors are also grateful to the Jet Propulsion Laboratory and the Minor Planet Center for providing astrometric data through the HORIZONS and MPC services, and to colleagues in the cometary physics community for constructive discussions that helped refine the thermophysical modeling framework. No external funding beyond institutional support was used for this study.
Data Availability
The observational data underlying this work were obtained from publicly available sources: the JPL HORIZONS ephemeris system (https://ssd.jpl.nasa.gov/horizons) and the Minor Planet Center database (https://minorplanetcenter.net). All code used to generate the thermophysical and Monte Carlo simulations is available as supplementary material. The figures and derived numerical results presented in this article can be reproduced directly from the scripts and data provided therein.
Appendix A Material Constants and Vapor Pressure Relations
@@
Table 2: Physical and thermodynamic constants used for volatile species in the model.
@@
Appendix B Model Implementation and Availability
@@
Appendix C Sensitivity to Rotation Period and Thermal Inertia

### v3

@@ Abstract
Interstellar comets , Non-gravitational acceleration , Outgassing dynamics , Thermophysical modeling , Cometary volatiles , CO sublimation , CO2 sublimation , Small-body dynamics
journal: Icarus
[…]
organization=Leiden Institute of Advanced Computer Science (LIACS), Leiden University,addressline=Gorlaeus Gebouw – BE-vleugel, Einsteinweg 55, city=Leiden, postcode=2333 CC, state=, country=The Netherlands
{highlights}
We model CO- and CO2-driven outgassing using a physically self-consistent thermophysical framework and real HORIZONS astrometry for 3I/ATLAS.
Minimum active-area fractions required to reproduce the observed non-gravitational acceleration are quantified across nucleus radii of 0.5–3 km.
CO retains strong sublimation efficiency at 1–2 AU, while NH3 and CH4 underproduce recoil except very close to the Sun.
A scalar lower-bound analysis shows that mixed CO–CO2 solutions yield mathematically minimal active fractions, though full directional solutions require larger areas.
Results demonstrate that the observed acceleration of 3I/ATLAS is consistent with ordinary volatile-driven activity, without invoking exotic physical mechanisms.
@@ 2 Observational Constraints
A non-gravitational solution was retrieved from HORIZONS on 15 January 2026. This solution reports statistically significant radial and transverse components, expressed in units of AU day-2. Conversion to SI units uses
$1~\mathrm{AU~day^{-2}}=2.004\times 10^{1}~\mathrm{m~s^{-2}}.$
The reported Marsden parameters are
@@
These correspond to radial and transverse accelerations of
$a_{1}=(2.7\pm 0.4)\times 10^{-5}~\mathrm{m~s^{-2}},\qquad a_{2}=(1.2\pm 0.4)\times 10^{-5}~\mathrm{m~s^{-2}},$
yielding a net non-gravitational acceleration magnitude of approximately
$a_{\mathrm{ng}}=(3.0\pm 0.8)\times 10^{-5}~\mathrm{m~s^{-2}}$
@@
To verify that the inclusion of non-gravitational terms improves the orbital fit, we compared the residuals of the gravitational-only and full HORIZONS solutions. The gravitational-only solution exhibits systematic residual trends aligned primarily with the heliocentric radial direction. In contrast, the solution including non-gravitational parameters removes these trends and reduces the residuals to the level of the astrometric uncertainties. This behavior is characteristic of recoil forces arising from anisotropic mass loss and is consistent with expectations for an active cometary nucleus.
The magnitude of the inferred acceleration lies toward the upper end of the range observed for active comets in the inner Solar System. Typical non-gravitational accelerations near 1 AU span approximately $10^{-9}$ to $10^{-5}~\mathrm{m~s^{-2}}$, depending on nucleus size, volatile composition, and activity geometry (Yeomans2004; Combi2011). For comparison, the interstellar object 1I/‘Oumuamua exhibited a non-gravitational acceleration of order $3\times 10^{-6}~\mathrm{m~s^{-2}}$ (Micheli2018). The value inferred for 3I/ATLAS is larger, consistent with its visibly active, cometary nature and with recoil forces driven by volatile sublimation at heliocentric distances of order 1–2 AU.
@@ 3.3 Material Parameters and Model Inputs
The thermophysical constants governing volatile sublimation are summarized in Table 1. These include molecular masses, latent heats, and empirical vapor-pressure coefficients derived from laboratory measurements of relevant ices. Bulk nucleus properties, including albedo, emissivity, and density, are treated separately and are chosen to be representative of low-albedo cometary nuclei. The bulk density is fixed at $500\,\mathrm{kg\,m^{-3}}$, consistent with in situ measurements of porous cometary material (Patzold2016).
The surface energy balance incorporates both radiative cooling and latent heat removal due to sublimation. Because the rotation state, obliquity, and active-area geometry of 3I/ATLAS are not constrained, the solar incidence factor $\mu_{0}$ is evaluated using a diurnal average that accounts for rotation and obliquity. This approach provides a physically motivated, geometry-agnostic estimate of the mean absorbed solar flux and avoids assumptions about specific spin states or localized subsolar activity.
Given these inputs, the model self-consistently solves for the equilibrium surface temperature, sublimation mass flux, and resulting recoil acceleration for each volatile species as a function of heliocentric distance.
Figure 1 shows the modeled diurnal-averaged non-gravitational acceleration produced by CO, CO2, NH3, and CH4 over the range $0.4\lesssim r_{H}\lesssim 3$ AU. The calculation adopts a representative jet-collimation factor $\eta=2.5$, a nucleus radius $R_{\mathrm{nuc}}=5.6$ km, bulk density $\rho=500$ kg m-3, and a total active fraction $f_{\mathrm{act}}=0.02$.
Under these assumptions, CO produces the strongest recoil acceleration across the full heliocentric distance range considered. CH4 yields accelerations that are systematically lower than those of CO but remain significant at small heliocentric distances. NH3 and CO2 generate substantially weaker accelerations, reflecting the rapid suppression of their sublimation rates once radiative cooling dominates the surface energy balance. The relative ordering and separation of the curves highlight the strong dependence of recoil efficiency on volatile thermodynamics and demonstrate that CO-driven sublimation is uniquely effective at producing sustained non-gravitational acceleration at heliocentric distances relevant to the 3I/ATLAS observations.
Figure 1: Diurnal-averaged non-gravitational acceleration as a function of heliocentric distance for CO, CO2, NH3, and CH4. Accelerations are computed from the full thermophysical energy-balance model using laboratory vapor-pressure relations, a diurnal-average illumination geometry, and a jet-collimation factor $\eta=2.5$. The example shown adopts a nucleus radius $R_{\mathrm{nuc}}=5.6$ km, bulk density $\rho=500$ kg m-3, and total active fraction $f_{\mathrm{act}}=0.02$. The curves illustrate the relative efficiency of different volatile species in producing recoil acceleration at heliocentric distances relevant to the 3I/ATLAS observations.
@@ 4.3 Comparison with Observations
Because the direction of the non-gravitational acceleration is poorly constrained for 3I/ATLAS, we compare the modeled accelerations to observations using only the magnitude of the inferred signal. The observed scalar acceleration $|\vec{a}_{\mathrm{obs}}|$ is derived from the radial and transverse Marsden parameters $A_{1}$ and $A_{2}$ obtained from the HORIZONS solution. For each assumed nucleus radius and volatile configuration, we determine the minimum total active fraction $f_{\mathrm{act,tot}}$ required to satisfy
@@
This scalar-matching approach yields a strict lower bound on the required active area, independent of the unknown orientation of the recoil force. Any physically consistent vector solution would require an equal or larger active fraction.
Figure 2 shows the resulting required active fractions for nucleus radii between 0.5 and 3 km, evaluated at the heliocentric distance $r_{H}=1.36$ AU. For all volatile compositions, the required active fraction increases monotonically with nucleus size, reflecting the scaling of sublimation-driven recoil with active surface area divided by nucleus mass.
Among the single-species models, CO requires the smallest active fraction at all radii, while CO2 and NH3 demand progressively larger active areas. In particular, NH3 requires implausibly large active fractions even for kilometer-scale nuclei, effectively ruling it out as a dominant driver of the observed non-gravitational acceleration. CO2 alone also requires large active fractions at modest nucleus sizes, consistent with its comparatively low sublimation efficiency at the inferred surface temperatures.
The mixed CO–CO2 (50/50) curve lies below the single-species cases and represents a formal lower bound obtained by scalar addition of species contributions. This curve does not correspond to a unique or physically realizable surface configuration, as it neglects constraints on the direction of the recoil force and the spatial coherence of active regions. Any configuration that simultaneously reproduces both the magnitude and direction of the observed acceleration would require an equal or larger total active fraction. The mixture curve should therefore be interpreted strictly as a lower bound on the required activity, not as a realistic description of the volatile distribution on the surface of 3I/ATLAS.
Figure 2: Required total active fraction as a function of nucleus radius for single-volatile and mixed-volatile sublimation models. For each radius, the scalar non-gravitational acceleration produced by the model is matched to the observed magnitude at $r_{H}=1.36$ AU. Single-species curves (CO, CO2, NH3) illustrate the increase in required active area with nucleus size due to the mass-to-area scaling of recoil acceleration. The CO–CO2 (50/50) curve represents a strict lower bound obtained by scalar combination of species contributions and does not constitute a physically unique solution; any directionally consistent model would require equal or larger active fractions.
@@ 5.4 Acceleration and Thermophysical Profiles
Figure 1 presents the modeled diurnal-averaged non-gravitational acceleration as a function of heliocentric distance. Among the volatiles considered, CO produces the largest recoil acceleration across the full range of $0.4\lesssim r_{H}\lesssim 3$ AU. CH4 yields systematically weaker accelerations that remain significant only at small heliocentric distances, while NH3 produces substantially smaller recoil throughout. CO2 contributes the weakest acceleration over the entire range, reflecting its limited sublimation efficiency under the modeled surface conditions.
Figure 2 illustrates how the active fraction required to reproduce the observed acceleration magnitude depends on nucleus size. As expected from the scaling of recoil force with active area divided by nucleus mass, smaller nuclei require lower active fractions, while larger nuclei demand progressively greater surface activity. Mixed CO–CO2 results represent formal lower bounds obtained from scalar combination of species contributions and do not correspond to physically unique or directionally consistent solutions.
The thermophysical origin of these trends is clarified in Figure 3, which shows the latent-to-radiative power ratio
$\Lambda(r_{H})=\frac{ZL}{\varepsilon\sigma T^{4}}$
as a function of heliocentric distance. This quantity provides a direct measure of whether surface energy balance is dominated by sublimation or by thermal radiation. CO and CH4 exhibit $\Lambda\gg 1$ over the full distance range considered, indicating strongly sublimation-dominated regimes. NH3 transitions more rapidly toward radiative dominance with increasing distance, while CO2 remains in a radiatively dominated regime ($\Lambda\ll 1$) throughout.
Taken together, Figures 13 demonstrate that efficient recoil acceleration at heliocentric distances relevant to the 3I/ATLAS observations requires volatiles whose sublimation remains energetically dominant under diurnal-averaged illumination. Among the species considered, CO uniquely satisfies this condition over the observed range, naturally explaining its dominance in the modeled non-gravitational acceleration.
Figure 3: Latent-to-radiative power ratio $\Lambda=(ZL)/(\varepsilon\sigma T^{4})$ as a function of heliocentric distance for CO, CO2, NH3, and CH4, computed using diurnal-averaged illumination and the full thermophysical energy-balance model. Values $\Lambda\gg 1$ correspond to sublimation-dominated surface energy balance, while $\Lambda\ll 1$ indicates radiative dominance. CO and CH4 remain sublimation-dominated across 0.40.433 AU, whereas CO2 is radiatively dominated throughout, limiting its ability to generate significant recoil acceleration.
@@ 7 Conclusions
8 Material Constants and Vapor Pressure Relations
@@
Table 1: Thermophysical constants for volatile species used in the sublimation model. Latent heats $L_{v}$ and molecular masses $\mu$ are taken from laboratory measurements. The vapor-pressure coefficients $A_{i}$ and $B_{i}$ define the empirical relation $\ln P=A_{i}-B_{i}/T$, with $P$ in pascals and $T$ in kelvin.
@@
9 Model Implementation and Availability
@@
10 Sensitivity to Rotation Period and Thermal Inertia